Do Chapter 13 End-of-Chapter Problems It is projected that approximately 580,000 veterans will take advantage of the GI Bill for the 21st Century. Boots to Books is a course for all veterans, current military members, and their family members, friends, and supporters. The goal of Boots to Books is to assist deployed, postdeployed, and veteran students in making a positive transition from military to civilian life or from deployment to postdeployment life, including the acquisition of college survival skills Post-traumatic stress disorder (PTSD) is quite common among combat veterans. Suppose that a researcher wants to revise the Boots to Books course irn order to reduce the effects of PTSD. He recruits a group of combat veterans and collects data on the length of time they were in the military (factor A) and the potential components of the revised course (community service, physical activity, and journaling: factor B). He then assigns them one of the three components. The study will evaluate which components of the program are the most effective for reducing PTSD symptoms The results of the hypothetical study are summarized in the following data matrix. Each cell reports the average (M), the total (T) and the sum of squares (SS) of the symptom score (X) on the Clinician-Administered PTSD Scale (CAPS) of 7 veterans. Veterans are grouped according to the number of years they served in the military Factor B: Program Component Community Physical Service Activity Journaling M 59.0 M57.7M53.9 T = 413 | T 404 | T = 377 T SS 22 SS 43.43 SS 2.86 ROW1,194 Factor A: 2X2 - 134,877 Years in Military 3-4 M 54.4 M56.9M57.7 T= 381 | T-398 | T-404 | TROW2-1,183 SS 43.43 SS 22.86 SS75.43 COL1 794 COL2- COL3 = 802 781 The researcher performs an analysis of variance (ANOVA) to test the hypothesis that the populations defined by the treatment combinations are equal The results are presented in the following ANOVA table ANOVA Table Source df MS Between treatments Factor A FactorB AXB interaction 143.8333 2.8810 16.0476 124.9048 2.8810 8.0238 62.4524 10.90 0.50 1.40

AXB interaction Within treatments Total 124.9048 206.2857 350.1190 62.4524 10.90 36 5.7300 41 Use the significance level α-.05 to complete the following conclusions. The critical F value for factor A (years in military) is 4.11. Therefore, the main effect due to factor A is The critical F value for factor B (program component) is 3.26. Therefore, the main effect due to factor B is The critical F value for the interaction of factors A and B is 3.26. Therefore, the effect due to the interaction of factors A and B is Given the results of the preceding analysis, what does the researcher conclude? O Veterans with different lengths of service had different PTSD symptom scores. The severity of PTSD symptoms was unrelated to program component either by itself or in conjunction with length of service O Neither length of service nor program component was related to the severity of PTSD symptoms when looked at alone. However, there was an O Neither length of service nor program component was related to the severity of PTSD symptoms, whether examined alone or in conjunction with O Veterans with different lengths of service had different PTSD symptom scores. By itself, the program that veterans took was not associated with interaction between the two factors such that severity differed by program component differently depending on length of service (and vice versa) the other factor differences in severity; however, length of service resulted in different degrees of PTSD symptoms depending on program. Measure the effect size for factor A (years in military), factor B (program component), and the interaction by computing (partial) eta squares. (Express the value of each eta square as a percent.) n2 for factor A (years in military) n2 for factor B (program component) η 2 for the interaction-
Answer #1

All the three F values are greater than the respective critical F values

Main effect due to factor A is not significant

Main effect due to factor B is not significant

Main effect due to interaction is not significant

Option C

eta^2 for factor A = 2.881/350.119 = 0.0082

eta^2 for factor B = 16.0476/350.119 = 0.0458

eta^2 for interaction = 124.9048/350.119 = 0.3567